OpenStudy (anonymous):

Write an equation that goes through the points (0,1) and (4,4)

OpenStudy (anonymous):

do you know the formula to findinga slope?

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

\[Slope (m)= \frac{ y2-y1 }{ x2-x1}\] \[= \frac{ 4-1 }{ 4-0 }\] \[= \frac{ 3 }{ 4 }\] Pick one of the coordinates and substitute into y=mx+b to find "b" y=mx+b \[4=\frac{ 3 }{ 4 }(4)+b\] \[4=3+b\] \[b= 4- 3\] \[b= 1\] Now the equation is \[y=\frac{ 3 }{ 4 }x+1\]